OpenStudy (anonymous):

solve the equation 8t-(5t+2)=16

OpenStudy (nikato):

Distribute the -(5t+2) And get -5t -2

OpenStudy (nikato):

So 8t-5t-2=16 Combine like terms on the left 3t-2=16 Add 2 to both sides 3t=18 Divide both sides by 3. t= 18/3 =?

OpenStudy (anonymous):

8t-5t-2=16 Simplify 8t-5t=3t 3t-2=16 Add 2 to both sides 3t=18 Divide by 3 on both sides 3t/3= 18/3 t=6